Two consecutive odd integers have a product of 99. Find all sets of integers that satisfy this description.
Answer by stanbon(48558)

Two consecutive odd integers have a product of 99. Find all sets of integers that satisfy this description.
-----------------
1st: 2x-1
2nd: 2x+1
------------------
Equation:
(2x-1)(2x+1) = 99
4x^2-1 = 99
4x^2 = 100
x^2 = 25
x = 5 or -5
-----------------
If x = 5,
2x-1 = 9
2x+1 = 11
-----------------
If x = -5,
2x-1 = -11
2x+1 = -9
=======================
Cheers,
Stan H.